Suppose that in a bushel of 100 apples there are 20 that have worms in them


and 15 that have bruises. Only those apples with neither worms nor bruises can


be sold. If there are 10 bruised apples that have worms in them, how many of


the 100 apples can be sold?

1
Expert's answer
2021-12-21T04:22:14-0500
"N(W\cup B)=N(W)+N(B)-N(W\cap B)"

"=20+15-10=25"

"100-25=75"

75 apples can be sold.